What is the natural domain of the function f(x,y)=√(x^2-y^2 )?
1
Expert's answer
2020-11-15T18:31:03-0500

Since "x^2-y^2\geq 0" , it follows that "x^2 \geq y^2" , that is, "|x| \geq |y|" . This domain is contained between lines "y=-x" and "y=x" if "x \geq 0", and between lines "y=x" and "y=-x" if "x<0" , including these lines.